ESEM is down for the weekend with a vacuum problem. I’ll fix it Monday
ESEM in Electron Optics
1:56 p.m. ESEM filament is back online and working fine. Ready to use.
12:15 p.m. The filament burned yesterday after. We put a new one in this morning but its not working just right. We are working on it now.
It’s too hot to touch today, so we’ll get it back online tomorrow morning.
There was a report of no EDS spectra on the Aztek system. On examination we found 2 user-correctable issues. First, the backscatter detector was blocking part of the x-ray signal from reaching the detector. It was not in the correct orientation. Then, and this is the biggest issue, the Process Time was set to 2 and it should be on 5. Process Time 2 provides narrower peaks (energy resolution) but reduces the deadtime (count rate). The SOP lists 5 as the default position. If you change it, please return to 5 when your work is complete.
The EDS works fine and is ready to use.
We’ve had a few ESEM problems but I think they are sorted out now. It’s okay to use the scope. Let me know if you see anything unusual.
Ready to use.
There is a problem with the ESEM software/computer. Don’t use it. We’ll get it back online tomorrow.
1:03 p.m.: There is a problem with the stage. I’ll check it and update you later this afternoon.
4:09 p.m.: Not sure what I did but its working now.
8:47 a.m.: The Oxford PC is down so EDS and EBSD can’t be used. I have a call into IT and I’ll update you as the repair is made.
10:31 a.m.: The Oxford pc is working now. Okay to use the scope.
The ESEM is back online. There is a bit of a problem with filament saturation but we are still working on that. The system is usable in the meantime.